By using your obtained surface area and its error S=3756.321(+or-)321,determine ,what will be a paint layer height, if the paint covering the cylinder has volume V=879(+or-)15.7,if the obtained height h=59.433 and radius r=8.77

Grade:12th pass

2 Answers

Arun
4 years ago
Dear student
 
Question is not clear. Please repost it with an attachment. I will be happy to help you.
Vikas TU
4 years ago
Dear student 
Surface area of cylinder = 2πr (r+h) 
Taking upper range , Surface area = 3756.321  + 321 = 4077.321
Upper unit of volume = 879 + 15.7 = 894.7
Volume = πr^2h 
Taking radius same  8.77
h = 894.7 / πr^2 
h = 3.702 unit 
Similarly apply for lower unit , in practical scenario , you should take maximum and calculate, 
Good Luck